OpenStudy (133):

What is the answer to 1+1x30+56-955=?

OpenStudy (anonymous):

-839

OpenStudy (anonymous):

-868

OpenStudy (anonymous):

1+1*30+56-955=-839

OpenStudy (mathlegend):

It would be nice to help the person out by doing the process along with the person. @Manny77 and @redrockett96 ... also you would know @redrockett96 by knowing PEMDAS... that your answer is incorrect.

OpenStudy (mathlegend):

@133 what needs to be done is you need to use PEMDAS. The 1st one that comes up based off of PEMDAS is multiplication right? 1+1x30+56-955

OpenStudy (anonymous):

1+1*30+56-955= 1+30+56-955= 31+56-955= 87-955= 868

OpenStudy (mathlegend):

Thank You @Manny77 ... you just forgot the negative sign this time. :)

OpenStudy (anonymous):

oh yea haha my bad its -868

OpenStudy (anonymous):

thanks for the medal ;)